VOLUME OF SOUND IN DECIBELS (dB) If the sound volume increases from 20 dB to 100 dB, then the sound volume increases by a factor of: Note: Remember that you can enter large numbers like 100000 as le5. Submit Answer Tries 0/3 If the sound volume decreases from 70 dB to 30 dB, then the sound volume decreases by a factor of: Submit Answer Tries 0/3
Answer #1

intensity level B = 10*log(I/Io)


for B1 = 20 dB = 10*log(I1/Io)


for B2 = 80 dB = 10*log(I2/Io)


B2 - B1 = 80 - 20


10*log(I2/Io) - 10*log(I1/Io) = 60


10*log(I2/I1) = 60


I2/I1 = 10^6 = 1e6

========================


intensity level B = 10*log(I/Io)


for B1 = 70 dB = 10*log(I1/Io)


for B2 = 30 dB = 10*log(I2/Io)


B2 - B1 = 30 - 70


10*log(I2/Io) - 10*log(I1/Io) = -40


10*log(I2/I1) = -40


I2/I1 = 10^-4 = 1e-4